OpenStudy (anonymous):

10 1/3 Multiplied by 1/9 in fractions.

OpenStudy (anonymous):

31/27

OpenStudy (anonymous):

10 1/3 = 31/3 31 1 31 so 10 1/3 * 1/9 = ------- * ----- = ----- 3 9 27

OpenStudy (anonymous):

you sure because i got 31/27 and it reduced to 3 4/9

OpenStudy (anonymous):

1 4/27 you mean.

OpenStudy (anonymous):

yes thanks a lot

OpenStudy (anonymous):

\[\frac{31}{27} = 1\frac{4}{27}\]